OpenStudy (anonymous):

Show that the product of z = x+yi and the expression z^-1 equals 1

OpenStudy (dan815):

okay so 1/z = z^-1

OpenStudy (dan815):

(x+yi) ----- (x+yi)

OpenStudy (dan815):

if you are uncomfortable with saying that is 1, multiply top and bottom by conjugate

OpenStudy (dan815):

(x+yi) (x-iy) x^2+y^2 -------------- = ---------- =1 (x+yi) (x-iy) x^2+y^2

OpenStudy (anonymous):

I understand, Thank you so much! It is simpler than it looks